Solve the folllowing system of linear equations
x-4y+2z=-1
2x+y-3z=10
-3x+12y-6z=3

Answer by Alan3354(69443)

x-4y+2z=-1
2x+y-3z=10
-3x+12y-6z=3
----------------
The 3rd eqn is the 1st multiplied by -3, so no solution is possible.